OpenStudy (anonymous):

can someone plz help with this question Solve. x2 – 10x = 5

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

8x=5 what are the answer chioces

OpenStudy (anonymous):

choices are: a. 5+-2sqrt5 b.5+-sqrt30 c.5+-sqrt15 d.5+-sqrt10

OpenStudy (anonymous):

i think it b

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

x^2-10x-5=0 Completing the square get: x^2-10x+25 = 5+25 (x-5)^2 = 30 x-5=sqrt30 or x-5=-sqrt30 x=5+sqrt30 or x=5-sqrt30